# Non-perturbative renormalization group for Higgs-like models in 4D

### v1

@@ III Projecting onto positive norm states and unitarity
$H^{\dagger}={\cal K}\,H\,{\cal K}^{\dagger},~{}~{}~{}{\cal K}^{\dagger}{\cal K}=1,~{}~{}~{}~{}~{}{\cal K}^{\dagger}=K,~{}~{}~{}~{}\Longrightarrow~{}~{}~{}{\cal K}^{2}=1.$ (23)
@@
In simple words, kets have an accompanying ${\cal K}$ but bras do not. In general, this new inner product has negative norm states. Below we will show that such states can be consistently projected out.
@@
$A^{\dagger_{\scriptscriptstyle k}}\equiv{\cal K}\,A^{\dagger}\,{\cal K}.$ (25)
@@
$H^{\dagger_{\scriptscriptstyle k}}=H.$ (26)
$A^{\dagger_{\scriptscriptstyle k}}$ is the proper conjugation based on the ${\cal K}$-metric in (24), namely
$\langle\langle\psi^{\prime}|A|\psi\rangle\rangle^{*}=\langle\langle\psi|A^{\dagger_{\scriptscriptstyle k}}|\psi^{\prime}\rangle\rangle,$ (27)
where * denotes ordinary complex conjugation. We define a pseudo-hermitian operator $A$ as one that satisfies $A^{\dagger_{\scriptscriptstyle k}}=A$. From (27), one concludes that pseudo-hermitian operators, in particular the hamiltonian H, has real eigenvalues. Any operator satisfying $A^{\dagger_{\scriptscriptstyle k}}=A$ in principle can correspond to an observable. One can easily establish that this pseudo-hermitian adjoint satisfies the usual rules, e.g.
$(AB)^{\dagger_{\scriptscriptstyle k}}$ $=$ $B^{\dagger_{\scriptscriptstyle k}}A^{\dagger_{\scriptscriptstyle k}}\,,$
$(aA+bB)^{\dagger_{\scriptscriptstyle k}}$ $=$ $a^{*}A^{\dagger_{\scriptscriptstyle k}}+b^{*}B^{\dagger_{\scriptscriptstyle k}}\,,$ (28)
@@
In general, the modified ${\cal K}$-metric $\langle\langle\psi^{\prime}|\psi\rangle\rangle$ introduces negative norm states which imply negative probabilities. The above statement of unitarity just states that the sum of these probabilities is maintained under time evolution, but one must still deal with negative norm states.
@@
which implies (38). The condition (39) implies $A^{\dagger}={\cal K}A^{\dagger}{\cal K}$. Thus $A^{\dagger_{\scriptscriptstyle k}}=A$ requires $A^{\dagger}=A$.
@@ IV Higher order RG beta-functions from the OPE
In this section we compute the beta functions to 3-loops for the fully anisotropic case $g_{1}\neq g_{2}\neq g_{3}$ based only on the OPE (15). We are more interested in the case where SU(2) is broken to U(1), where $g_{1}=g_{2}\neq g_{3}$ considered in the next section, however it’s easier to work out the beta functions for the fully anisotropic case first and then set $g_{1}=g_{2}$. Renormalization group beta-functions are generally prescription, or scheme, dependent. However for marginal perturbations, the one and two loop contributions are universal. [^5] To see this, consider a single coupling $g$ with $\beta_{g}=b_{2}g^{2}+b_{3}g^{3}+\ldots$ where $b_{2,3}$ are coefficients. The prescription dependence corresponds to a redefinition of the coupling $g^{\prime}=g^{\prime}(g)$. Let $g^{\prime}=g+cg^{2}+\ldots$. One easily sees that $\beta^{\prime}(g^{\prime})=b_{2}g^{\prime 2}+b_{3}g^{\prime 3}+\ldots.$ The implicit prescription of this article is based on the OPE (15), and doesn’t rely on any epsilon-expansion, nor Feynman diagrams for scalar fields. [^6] We do not address whether this OPE prescription can be shown to be equivalent to Feynman diagram perturbation theory and dimensional regularization. Referring to the title of Wilson-Fisher’s paper WilsonFisher , we work in precisely $4$ dimensions, not $3.99$! We first review the 1-loop result in ALDolls , then describe how to extend this to higher orders. The starting point is the perturbative expansion about the conformal field theory for $\langle{\bf{X}}\rangle$ where ${\bf{X}}$ is an arbitrary operator (it can in fact to be taken to be the identity):
@@ VI.2 Massless Flows between two non-trivial fixed points for imaginary coupling
Massive theories typically terminate at an empty fixed point, since at low energies the massive particles decouple. On the other hand, flows that end at a non-trivial CFT are expected to be massless, since some massless degrees of freedom survive the flow. There are no such flows in our models if $g_{1}$ and $g_{3}$ are real. Well, since we are already considering non-unitary theories, consider $g_{1}$ to be an imaginary coupling:
@@
Figure 11: Contour Plot of the RG invariant $Q$ for $g_{1}$ imaginary.
@@ VI.3 Some special rational points in the 4D CFT’s
We have already seen that the point $g_{3}=4/3$ is special since it is at the boundary of the massless flows where $\Gamma_{\scriptscriptstyle{\rm IR}}=\infty$. One can surmise other potentially interesting special points as follows. When $\Gamma_{\scriptscriptstyle{\rm IR}}=\mathfrak{D}$ where $\mathfrak{D}$ is an integer $\mathfrak{D}\geq 4$, then the IR theory has some likely interpretations. For instance, if $\mathfrak{D}=8$, this could represent massless flows that arrive to the IR via the operator $T_{\mu\nu}T^{\mu\nu}$, or $(T_{\mu}^{\mu})^{2}$ where $T_{\mu\nu}$ is the dimension 4 energy-momentum tensor. If this were the case, then this is analogous to $T\overline{T}$ perturbations of 2D CFT SmirnovZamo . In this case $\Gamma_{\scriptscriptstyle{\rm UV}}=16/5$ which is rational and corresponds to $g_{3}=1$, which is in the range (79). If one repeats this argument for the same beta function in 2 spacetime dimensions based on (97), one finds $\Gamma_{\scriptscriptstyle{\rm UV}}=4/3$. This turns out to be the scaling dimension of the perturbation for the minimal ${\cal N}=2$ supersymmetric model at Virasoro central charge $c=1$ BLsusy , which is a special point of the sine-Gordon model at coupling $b^{2}=4/3$ in (93). Thus the 2D analog of this flow is such that it preserves the central charge $c$, but the ${\cal N}=2$ supersymmetry is broken in the flow and the remaining massless fields are goldstinos. [^9] To our knowledge this observation is absent from the literature. This flow is similar to massless flows that arrive in the IR to the CFT of a Majorana fermion via the dimension 4 operator $T\overline{T}$ Alyosha , where the UV central charge is $7/10$. It was shown in AhnLeClair that the only UV completions of flows that end at the Majorana description of the Ising model have UV central charge $c=7/10$ and $c=3/2$, both of which have ${\cal N}=1$ supersymmetry. [^10] There are other flows that arrive to the Ising model via $T\overline{T}$ based on a different spectrum related to the $E_{8}$ Lie algebra AhnLeClair .
@@
In this limit $\Gamma_{\scriptscriptstyle{\rm UV}}=3$. This implies that the coupling $g_{1}$ has dimension $1$ in the UV, i.e. it is a like a mass coupling. One can argue that this special point is a free field theory, since there is nothing to flow to if $\Gamma_{\scriptscriptstyle{\rm IR}}>\infty$, thus there are effectively no flows. Much more work would be needed to establish this. However, in support of this idea, let us point out that for the models with the same beta function in 2D reviewed in the Appendix where (77) is replaced with (97), there the analog of this special point is $\Gamma_{\scriptscriptstyle{\rm UV}}=1$, which is known to be a mass term for a free Dirac fermion in 2D. [^11] Since relativistic fermion fields $\psi$ have dimension $3/2$ in 4D, this suggests that $g_{3}=4/3$ could also be a free fermion point, where the perturbation is $m\overline{\psi}\psi$ with $m$ identified by $g_{1}$. However lacking a formalism of bosonization in 4D, this would be difficult to establish at this stage, and probably unlikely.
@@ VI.4 Cyclic RG flows
Cyclic flows have previously been considered as rare, or just curiosities, because of various c-theorems and other considerationsctheorem ; CardyCtheorem ; Osborn ; JackOsborn ; Schwimmer ; Komargodski ; Polchinski . The latter studies often utilize quantum fields coupled to gravity since the focus is on the conformal, or trace anomaly of the stress-energy tensor. All of these works assume the theory is unitary. To address this further, first and foremost, our model certainly exists by construction, and as we showed, it has a regime with massless flows between conformal fixed points, which are more conventional flows. Although the CFT’s are non-unitary, these massless flows still fit in the paradigm that all flows begin and end at conformal fixed points. One should mention that in recent studies of massless flows between non-unitary minimal CFT’s in 2D, such flows are still consistent with a c-theorem Ravanini ; MussardoLG ; Tanaka ; KlebanovMM ; Negro . Returning to our model, if one is forced to understand the flows for all regimes of the couplings $g_{1},g_{3}$, then a cyclic regime where $Q$ is positive is an inescapable conclusion. At small coupling, this is just above the SU(2) symmetric diagonals. Secondly, for the 2D models with beta functions that are identical to those considered here and reviewed in the Appendix, an exact S-matrix was even proposed LeClairSierra , further justifying that these models are well-defined.
@@ VII Conclusions
It would be desirable to have a more fundamental understanding of the ${\cal K}$-insertions for our modified metric on the Hilbert space, since in this article it was just introduced in order to obtain the interesting OPE (15) since this structure leads to interesting beta-functions. Very recent work of Witten is suggestive of a such an understanding, where he argues that bras and kets can differ by an operator, which is analogous to our ${\cal K}$-inner product where kets have an operator ${\cal K}$ insertion, whereas bras do not.
@@ VIII Acknowledgements
We wish to thank Luis Alvarez-Gaumé, Denis Bernard, Igor Klebanov, Michael Peskin, and Germán Sierra for discussions.
@@ IX Appendix: Comparison with 2D current-current perturbations
$\bullet$ When $g_{1}$ is imaginary, i.e. $g_{1}\to ig_{1}$, then this is the so-called imaginary sine-Gordon model Fendley1 ; Fendley2 , which has massless flows between two fixed points on the critical line $g_{1}=0$. These flows do not change the Virasoro central charge $c=1$, and the UV and IR fixed points are those of a free boson with different compactification radius. That the central charge does not change already violates the c-theorem ctheorem , presumably because of the non-unitarity. Repeating the arguments of Section VI based on the beta functions, the relation between the anomalous dimensions in the UV and IR is
@@
These flows only exist for $\Gamma_{\scriptscriptstyle{\rm UV}}>1$, where $\Gamma_{\scriptscriptstyle{\rm UV}}=1$ corresponds to the free-fermion point of the sine-Gordon model. This model is integrable and the relation (97) agrees exactly with the thermodynamic Bethe-ansatz analysis in Fendley2 , and this provides a non-trivial check of the validity of our beta functions to all orders.
Figure 13: A rough sketch of non-perturbative flows. A proper rendition of these flows based on the RG invariant $Q$ is shown in Figure 10. This figure borrowed from BLflow : sG, shG, and KT stand for sine-Gordon, sinh-Gordon, and Kosterlitz-Thouless.

### v2

@@ III Projecting onto positive norm states and unitarity
$H^{\dagger}={\cal K}\,H\,{\cal K}^{\dagger},~{}~{}~{}{\cal K}^{\dagger}{\cal K}=1,~{}~{}~{}~{}~{}{\cal K}^{\dagger}={\cal K},~{}~{}~{}~{}\Longrightarrow~{}~{}~{}{\cal K}^{2}=1.$ (26)
@@
In simple words, kets have an accompanying ${\cal K}$ but bras do not, or visa-versa since ${\cal K}^{\dagger}={\cal K}$. In general, this new inner product has negative norm states. Below we will show that such states can be consistently projected out.
@@
$A^{\dagger_{\scriptscriptstyle\kappa}}\equiv{\cal K}\,A^{\dagger}\,{\cal K}.$ (28)
@@
$H^{\dagger_{\scriptscriptstyle\kappa}}=H.$ (29)
$A^{\dagger_{\scriptscriptstyle\kappa}}$ is the proper conjugation based on the ${\cal K}$-metric in (27), namely
$\langle\!\langle\psi^{\prime}|A|\psi\rangle\!\rangle^{*}=\langle\!\langle\psi|A^{\dagger_{\scriptscriptstyle\kappa}}|\psi^{\prime}\rangle\!\rangle,$ (30)
where * denotes ordinary complex conjugation. We define a pseudo-hermitian operator $A$ as one that satisfies $A^{\dagger_{\scriptscriptstyle\kappa}}=A$. From (30), one concludes that pseudo-hermitian operators, in particular the hamiltonian H, has real eigenvalues. Any operator satisfying $A^{\dagger_{\scriptscriptstyle\kappa}}=A$ in principle can correspond to an observable. One can easily establish that this pseudo-hermitian adjoint satisfies the usual rules, e.g.
$(AB)^{\dagger_{\scriptscriptstyle\kappa}}$ $=$ $B^{\dagger_{\scriptscriptstyle\kappa}}A^{\dagger_{\scriptscriptstyle\kappa}}\,,$
$(aA+bB)^{\dagger_{\scriptscriptstyle\kappa}}$ $=$ $a^{*}A^{\dagger_{\scriptscriptstyle\kappa}}+b^{*}B^{\dagger_{\scriptscriptstyle\kappa}}\,,$ (31)
@@
In general, the modified ${\cal K}$-metric $\langle\!\langle\psi^{\prime}|\psi\rangle\!\rangle$ introduces negative norm states which imply negative probabilities. The above statement of unitarity just states that the sum of these probabilities is maintained under time evolution, but one must still deal with negative norm states which implies negative probabilities.
@@
which implies (41). The condition (42) implies $A^{\dagger}={\cal K}A^{\dagger}{\cal K}$. Thus $A^{\dagger_{\scriptscriptstyle\kappa}}=A$ requires $A^{\dagger}=A$.
@@
Above we interpreted the operators $a_{\pm}^{\dagger}$ as creating ${\cal P}{\cal T}$ even/odd states. Thus projecting onto ${\cal H}_{+}$ breaks ${\cal P}{\cal T}$ symmetry.
@@ IV Higher order RG beta-functions from the OPE
In this section we compute the beta functions to 3-loops for the fully anisotropic case $g_{1}\neq g_{2}\neq g_{3}$ based only on the OPE (18). We are more interested in the case where SU(2) is broken to U(1), where $g_{1}=g_{2}\neq g_{3}$ considered in the next section, however it’s easier to work out the beta functions for the fully anisotropic case first and then set $g_{1}=g_{2}$. Renormalization group beta-functions are generally prescription, or scheme, dependent. However for marginal perturbations, the one and two loop contributions are universal. [^5] To see this, consider a single coupling $g$ with $\beta_{g}=b_{2}g^{2}+b_{3}g^{3}+\ldots$ where $b_{2,3}$ are coefficients. The prescription dependence corresponds to a redefinition of the coupling $g^{\prime}=g^{\prime}(g)$. Let $g^{\prime}=g+cg^{2}+\ldots$. One easily sees that $\beta^{\prime}(g^{\prime})=b_{2}g^{\prime 2}+b_{3}g^{\prime 3}+\ldots.$ The implicit prescription of this article is based on the OPE (18), and doesn’t rely on any epsilon-expansion, nor Feynman diagrams for scalar fields. [^6] We do not address whether this OPE prescription can be shown to be equivalent to Feynman diagram perturbation theory and dimensional regularization. Referring to the title of Wilson-Fisher’s paper WilsonFisher , we work in precisely $4$ dimensions, not $3.99$!
We first review the 1-loop result in ALDolls , then describe how to extend this to higher orders. The starting point is the perturbative expansion about the conformal field theory for $\langle{\bf{X}}\rangle$ where ${\bf{X}}$ is an arbitrary operator (it can in fact to be taken to be the identity):
@@ VI.2 Massless Flows between two non-trivial fixed points for imaginary coupling
Massive theories typically terminate at an empty fixed point, since at low energies the massive particles decouple. On the other hand, flows that end in the IR at a non-trivial CFT are expected to be massless, since some massless degrees of freedom survive the flow. There are no such flows in our models if $g_{1}$ and $g_{3}$ are real. Well, since we are already considering non-unitary theories, consider $g_{1}$ to be an imaginary coupling:
@@
Figure 11: Contour Plot of the RG invariant $Q$ for $g_{1}$ imaginary. These flow contours begin and end along the critical $g_{3}$ line, and do not pass through it as the above contours do.
@@ VI.3 Some special rational points in the 4D CFT’s
We have already seen that the point $g_{3}=4/3$ is special since it is at the boundary of the massless flows where $\Gamma_{\scriptscriptstyle{\rm IR}}=\infty$. One can surmise other potentially interesting special points as follows. When $\Gamma_{\scriptscriptstyle{\rm IR}}=\mathfrak{D}$ where $\mathfrak{D}$ is an integer $\mathfrak{D}\geq 4$, then the IR theory has some likely interpretations. For instance, if $\mathfrak{D}=8$, this could represent massless flows that arrive to the IR via the operator $T_{\mu\nu}T^{\mu\nu}$, or $(T_{\mu}^{\mu})^{2}$ where $T_{\mu\nu}$ is the dimension 4 energy-momentum tensor. If this were the case, then this is analogous to $T\overline{T}$ perturbations of 2D CFT SmirnovZamo . In this case $\Gamma_{\scriptscriptstyle{\rm UV}}=16/5$ which is rational and corresponds to $g_{3}=1$, which is in the range (82). If one repeats this argument for the same beta function in 2 spacetime dimensions based on (100), one finds $\Gamma_{\scriptscriptstyle{\rm UV}}=4/3$. This turns out to be the scaling dimension of the perturbation for the minimal ${\cal N}=2$ supersymmetric model at Virasoro central charge $c=1$ BLsusy , which is a special point of the sine-Gordon model at coupling $b^{2}=4/3$ in (96). Thus the 2D analog of this flow is such that it preserves the central charge $c$, but the ${\cal N}=2$ supersymmetry is broken in the flow and the remaining massless fields are goldstinos. [^9] To our knowledge this observation is absent from the literature. This flow is similar to massless flows that arrive in the IR to the CFT of a Majorana fermion via the dimension 4 operator $T\overline{T}$ Alyosha , where the UV central charge is $7/10$. It was shown in AhnLeClair that the only UV completions of flows that end at the Majorana description of the Ising model have UV central charge $c=7/10$ and $c=3/2$, both of which have ${\cal N}=1$ supersymmetry. [^10] There are other flows that arrive to the Ising model via $T\overline{T}$ based on a different spectrum related to the $E_{8}$ Lie algebra AhnLeClair , indicating multi-critical behavior. In the notation in Tanaka , this is a flow between $c<1$ minimal models $M(11,12)\to M(3,4)$, with central charge $c_{\scriptscriptstyle{\rm UV}}=21/22$, $c_{\scriptscriptstyle{\rm IR}}=1/2$. This flow is not included in Tanaka since there the focus on the generic SU(2) quantum group modular fusion category of Verlinde lines, whereas our $E_{8}$ based flow should be understood from the $E_{8}$ quantum group fusion category.
@@
In this limit $\Gamma_{\scriptscriptstyle{\rm UV}}=3$. This implies that the coupling $g_{1}$ has dimension $1$ in the UV, i.e. it is a like a mass coupling. One can argue that this special point is a free field theory, since there is nothing to flow to if $\Gamma_{\scriptscriptstyle{\rm IR}}>\infty$, thus there are effectively no flows. Much more work would be needed to establish this. However, in support of this idea, let us point out that for the models with the same beta function in 2D reviewed in the Appendix where (80) is replaced with (100), there the analog of this special point is $\Gamma_{\scriptscriptstyle{\rm UV}}=1$, which is known to be a mass term for a free Dirac fermion in 2D. [^11] Since relativistic fermion fields $\psi$ have dimension $3/2$ in 4D, this suggests that $g_{3}=4/3$ could also be a free fermion point, where the perturbation is $m\overline{\psi}\psi$ with $m$ identified by $g_{1}$. However lacking a formalism of bosonization in 4D, this would be difficult to establish at this stage.
@@ VI.4 Cyclic RG flows
Cyclic flows have previously been considered as rare, or just curiosities, because of various c-theorems and other considerationsctheorem ; CardyCtheorem ; Osborn ; JackOsborn ; Schwimmer ; Komargodski ; Polchinski . The latter studies often utilize quantum fields coupled to gravity since the focus is on the conformal, or trace anomaly of the stress-energy tensor. All of these works assume the theory is unitary. To address this further, first and foremost, our model certainly exists by construction, and as we showed, it has a regime with massless flows between conformal fixed points, which are more conventional flows. Although the CFT’s are non-unitary, these massless flows still fit in the paradigm that all flows begin and end at conformal fixed points. One should mention that in recent studies of massless flows between non-unitary minimal CFT’s in 2D, such flows are still consistent with a c-theorem Ravanini ; MussardoLG ; Tanaka ; KlebanovMM ; Negro , where here $c=c_{{\rm eff}}$, where $c_{{\rm eff}}$ is the effective central charge that appears in the thermodynamic Bethe ansatz. Returning to our model, if one is forced to understand the flows for all regimes of the couplings $g_{1},g_{3}$, then a cyclic regime where $Q$ is positive is an inescapable conclusion. At small coupling, this is just above the SU(2) symmetric diagonals. Secondly, for the 2D models with beta functions that are identical to those considered here and reviewed in the Appendix, an exact S-matrix was even proposed LeClairSierra , further justifying that these models are well-defined.
@@ VII Conclusions
It would be clearly desirable to have a more fundamental understanding of the ${\cal K}$-insertions for our modified metric on the Hilbert space, since in this article it was just introduced in order to obtain the interesting OPE (18) since this structure leads to interesting beta-functions. Based on our results on massless flows above, by analogy with the classification of massless flows between non-unitary CFT;s in 2D Tanaka , we suggest that ${\cal K}$ can be interpreted as a topological defect, analogous to Verlinde lines in 2D CFT. In the 4D context, these should be codimension 2 defects, namely two dimensional surfaces. It is well accepted that such defects can modify the critical behavior in significant ways. The massless flows we found along the critical line $g_{1}=0$ should then be viewed as constrained by the U(1) fusion category. From this viewpoint, the line of fixed point we found are analogous to the radius of compactification of a free boson in 2D, which also has a duality. The flows in Tanaka are based on a study of the $SU(2)_{q}$ quantum group modular fusion category where $q$ is a root of unity, which is known to be a reduction of the U(1) category from various viewpoints. Very recent work of Witten is also suggestive of a such an understanding, where he argues that bras and kets can differ by an operator, which is analogous to our ${\cal K}$-inner product where kets have an operator ${\cal K}$ insertion, whereas bras do not.
There is some evidence for an underlying ${\rm SL}(2,\mathbb{Z})$ structure in our model. Above we showed that the higher order beta functions have a $g\to 1/g$ duality, like an S-duality $\tau\to-1/\tau$. T-duality, $\tau\to\tau+1$ typically corresponds to a periodicity due to a $\vartheta$-angle, as in ${\cal N}=4$ supersymmetric Yang-Mills theory. The cyclicity of the RG required us to identify $g_{3}=\pm\infty$, endowing the 2-parameter coupling constant space $(g_{1},g_{3})$ with the topology of a cylinder, which is a form of periodicity. This by itself does not establish an ${\rm SL}(2,\mathbb{Z})$ symmetry, since we have not precisely identified the elliptic modulus $\tau$ as a function of the 2 couplings. The role of ${\rm SL}(2,\mathbb{Z})$ is perhaps clearer for the fully anisotropic case with the beta functions (53) since integration of these 1-loop beta functions already leads to elliptic functions LeClairSierra . Furthermore the 2D models, with identical beta functions to our models, were proposed to have the S-matrix of Zamolodchikov Zelip built out of elliptic functions Elliptic . This S-matrix is characterized by a $\mathbb{Z}_{4}$ symmetry which is a subgroup of ${\rm SL}(2,\mathbb{Z})$, generated by the matrix $A=-i\sigma^{2}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$. Furthermore, it is known that this S-matrix transforms covariantly under ${\rm SL}(2,\mathbb{Z})$, meaning the S-matrix transforms with phase factors or gauge transformations that preserve the structure of the Yang-Baxter equation Baxter ; Belavin . In Elliptic , for the SU(2) broken to U(1) model, which is the cyclic sine-Gordon model reviewed in the Appendix, the latter can be obtained from certain limits of Zamolodchikov’s elliptic S-matrix, and this is perhaps a way of understanding what is left of the ${\rm SL}(2,\mathbb{Z})$ structure for our two parameter model.
@@ VIII Acknowledgements
We wish to thank Luis Alvarez-Gaumé, Denis Bernard, Igor Klebanov, Michael Peskin, Germán Sierra and Takahiro Tanaka for discussions.
@@ IX Appendix: Comparison with 2D current-current perturbations
$\bullet$ When $g_{1}$ is imaginary, i.e. $g_{1}\to ig_{1}$, then this is the so-called imaginary sine-Gordon model Fendley1 ; Fendley2 , which has massless flows between two fixed points on the critical line $g_{1}=0$. These flows do not change the Virasoro central charge $c=1$, and the UV and IR fixed points are those of a free boson with different compactification radius. That the central charge does not change already violates the c-theorem ctheorem , presumably because of the non-unitarity. Quantum group reductions of these flows describe flows between unitary minimal models with $c<1$. Repeating the arguments of Section VI based on the beta functions, the relation between the anomalous dimensions in the UV and IR is
@@
These flows only exist for 1¡$\Gamma_{\scriptscriptstyle{\rm UV}}<2$, where $\Gamma_{\scriptscriptstyle{\rm UV}}=1$ corresponds to the free-fermion point of the sine-Gordon model. This model is integrable and the relation (100) agrees exactly with the thermodynamic Bethe-ansatz analysis in Fendley2 , and this provides a non-trivial check of the validity of our beta functions to all orders.
Figure 13: A rough sketch of non-perturbative flows for $g_{1}>0$. Flows for $g_{1}<0$ are a mirror reflection. A proper rendition of these flows based on the RG invariant $Q$ is shown in Figure 10. This figure borrowed from BLflow : sG, shG, and KT stand for sine-Gordon, sinh-Gordon, and Kosterlitz-Thouless.